Find the solution set of 18 - |2y-6| > 45
Find the values of k so that the line Y = KX + 10 will form a triangle of area 60 with the coordinate axis. Give your answers in simplified fraction form.
The female born in the later part of 20th century will be x years old in the year x squared. In what year was she born? (x is in integers)

Answer by stanbon(75887)

Find the solution set of 18 - |2y-6| > 45
----
-|2y-6| > 27
|2y-6| < -27
No solutions as an absolute value cannot be negative.
===========================================================
Find the values of k so that the line Y = KX + 10 will form a triangle of area 60 with the coordinate axis. Give your answers in simplified fraction form.
---
The y-intercept is (0,10)
Plot the point (0,10).
Note that the height of the triangle is 10.
So, if the area is 60,
Area = (1/2)*height*base
60 = (1/2)10*base
base = 12
So the x-intercept is 12
Plot (12,0)
---
Solve 0 = k*12 +10
---
k = -5/6
==========================================
The female born in the later part of 20th century will be x years old in the year x squared. In what year was she born? (x is in integers)
----
I'll leave that to you.
cheers,
Stan H.